• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name A. Camsonne

    Log entry time 08:44:54 on May 3,2008

    Entry number 236654

    keyword=Owl shift summary

    SL/TO : A. Camsonne
    3rd : K. Chirapatpimol


    0:00 taking over from swing shift no beam
    0:27 MCC says RF problem is fixed and expect beam back in about 30 minutes
    1:11 moving to empty target
    1:31 tune is good LD2 target in position
    Take calibration data run 4633 and Run 4634
    2:00 Moving to loop 1
    Take calibration data Run 4635 and 4636

    3:01 Bigbite set to 200 A, Left HRS Sieve in

    3:17 increased current from 4 to 5 uA

    4:41 Bigbite set to 500 A

    5:24 Big Bite set to 600 A

    5:58 BigBite HV trip

    6:08 BigBite current to 390 A

    6:38 Left HRS Sieve out

    6:56 starting kinematic change and cycling quads

    7:13 cycling done and magnets set

    7:20 Target moved to Ta target

    7:24 Angle set to 16.5

    7:28 Start data taking Kin B setting, 4 uA

    7:45 moving to C target

    8:07 Set Left HRS to 1.1338 GeV/c



    LD2 runs
    Run
    4633 500 K
    20850 76K
    4634 500 K

    LH2 runs
    4635 500 K run taken while changing fan speed and coolant
    4636 500 K run taken while changing fan speed and coolant
    20851 82K

    LH2 runs with Left HRS Sieve in
    4637 400 K BigBite current at 200 A
    20852 152K
    4638 200 K BigBite current at 500 A
    20853
    4639 141K BigBite Current at 600 A
    20854 54K
    4640 30K BigBite Current at 600 A
    20855 12K
    4641 81K BigBite Current at 390 A
    20856 31K

    LH2 run sieve out
    4642 200 K
    4643 200 K
    20857 29K

    Kin B runs
    Ta run
    4644 529 K
    20058 15K

    C run
    4645 721 K need to check offline dead time had a dead time of 45% at 1 KHz of trigger
    20059 13K